OpenStudy (anonymous):

Given that P(A)=0.4, P(b)=0.5,and (P and B)=0.20, determine P(A|B)

OpenStudy (anonymous):

\[P(A|B)=\frac{P(A\cap B)}{P(B)}=\frac{.2}{.5}=\frac{2}{5}\]